OpenStudy (anonymous):

simplify:

OpenStudy (anonymous):

\[Y=\frac{27s-9}{(9s^{2}-6s+1)}\]

OpenStudy (anonymous):

i get \[Y=\frac{9(3s-1)}{(s-\frac{1}{3})(s-\frac{1}{3})}\]

OpenStudy (anonymous):

but i dont think its in the simplified version

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

I get y = 9 /3s -1

OpenStudy (anonymous):

kay ._. cant see the relationship at first.. LOL brain clogged

OpenStudy (lgbasallote):

ahh you know the numerator...but \(\Large (\frac{1}{3})^{2}\) is not equal to 1 :P

OpenStudy (anonymous):

so it also equals to \[\frac{81}{3s-1}\]?

OpenStudy (lgbasallote):

81? no...remember you have \(\Large \frac{9(3s-1)}{(3s-1)(3s-1)}\)

OpenStudy (anonymous):

because if i factor another 3 out from the numerator, i can cancel 1 term of (s-1/3) frm top and btm

OpenStudy (anonymous):

omg == brain clogged again

OpenStudy (lgbasallote):

\[\huge 9s^{2} - 6s + 1 \ne (s - \frac{1}{3})(s - \frac{1}{3})\]

OpenStudy (anonymous):

thr's a 9 outside of (s-1/3)(s-1/3)?

OpenStudy (lgbasallote):

wait...let's start from the beginning..there seems to be some confusion..

OpenStudy (anonymous):

laplace derivative clogging my brain ...

OpenStudy (lgbasallote):

first factor out the numerator.. (27s - 9)..we take 9 out..9(3s - 1) \(\Large \mathsf{good?}\) ..wait is this laplace derivative?

OpenStudy (anonymous):

er is part of it, i need to simplify it b4 applying the inverse

OpenStudy (anonymous):

i think i know where i went wrong

OpenStudy (anonymous):

9s^2 -6s +1 is (3s-1)^2 == i wrote it as (s-1/3)^2

OpenStudy (lgbasallote):

yes! right :DDD